# Johannes Huemer [^2] These authors contributed equally to this work. Markus Murschitz33footnotemark: 3 Matthias Schörghuber Lukas Reisinger Thomas Kadiofsky Christoph Weidinger Mario Niedermeyer Benedikt Widy Marcel Zeilinger Csaba Beleznai Tobias Glück Andreas Kugi Patrik Zips44footnotemark: 4

### v1

@@ 4.4 Loading Edge Detection
Outlier Removal: Outliers are removed by discarding points that lack a sufficient number of neighbors within a specified radius.
@@
Figure 11: A LiDAR point cloud (colored by reflectance) captures the scene, including the forklift forks (bottom right). The lift is positioned at an estimated loading height to optimize the platform’s visibility. The resulting detected loading edge reference frame (RGB frame) and the target slots (red boxes) defined within this frame are depicted as overlay.
@@ 4.5 Obstacle Detection
(b) Birds-eye view: potential obstacles (green) are tracked to assess movement but are only considered as future collision (red) if they intersect with the path volume (cyan)
@@
Figure 12: Depiction of the obstacle detection system’s geometric relations and output. Definitive collision objects (red) that trigger emergency braking overlap with the expected path volume (cyan) and tracked potential obstacles are outside (green).
@@
For safe operation, the forklift must detect obstacles quickly to avoid collisions with both static obstacles (trees, walls, etc.) and moving obstacles (excavators, pedestrians, etc.) which may approach the currently planned drive path. The implemented approach is a fast-reacting method in the sensor frame complementary to and independent of the mapping approach presented in Section 4.2. It is applied only in the forward driving direction, where the risk of colliding with moving obstacles is higher, whereas movements in fork direction, as discussed in Section 3.3.2, are slower and follow much shorter paths. Based on the selected safety concept, when an obstacle is detected, the forklift is brought to a safe state (standstill) to prevent collisions, rather than re-planning its path to navigate around the obstacle. Once an emergency stop is triggered, the forklift remains stationary until it receives operator input to resume the task.
@@
For the assessment of the risk of collision with moving obstacles, we utilize a multi-object detection and tracking system, originally based on the works [72] and [73], which further became the basis for an endurance-tested tramway braking assistance system in commercial use [^5] https://www.alstom.com/solutions/rolling-stock/citadis-light-rail-designed-reflect-your-citys-unique-identity (accessed 2024-12-11). It operates in a tracking-by-detection paradigm. It can detect arbitrary objects, track them over time, and assess if a collision will occur based on the predicted trajectories of the objects and the projected path of the forklift. The detection step is based on generic 3D clustering, hence it does not depend on predefined object classes such as person or vehicle.
@@
To detect obstacles, the system analyzes point clouds from the Ouster LiDAR mounted on the front part of the forklift. Both LiDAR and a custom stereo camera setup, see Figure 3c, have been successfully used, but the wider field of view of the LiDAR system makes it the safer choice for obstacle detection, see Figure 12. Points above an assumed ground plane, cf. Figure 12(a), are analyzed for potential obstacles by grouping them into separate distinct objects using Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [74]. The clusters represent potential obstacles (indicated as red and green boxes in Figure 12) but only if they overlap with the expected vehicle’s path and are considered potential collision objects (red) that trigger an emergency brake. The volume of the monitored path, cyan in Figure 12, is defined by extruding the vehicle footprint (plus an additional clearance buffer) along a simplified vehicle path. Obstacle bookkeeping is implemented in a manner similar to pallet bookkeeping, as described in Section 4.1. However, it is enhanced with a Kalman filter based on a constant velocity model to accommodate movement. Then each object track is checked for a possible collision, which can cause a collision warning based on the predicted movement of the object, the velocity of the forklift, the characteristics of the forklift deceleration, and a defined minimum safety distance.
@@ 5 Planning and Control
ADAPT performs two primary types of actions to carry out its tasks effectively: (1) collision-free Navigation, where the forklift moves from its current position to a target location within a predefined operational area, either while carrying a pallet or traveling unloaded, and (2) precise Manipulation, which involves the accurate pickup and placement of pallets in specified locations (slots), such as on a truck or within a designated area on the ground (multiple slots on a truck are shown in Figure 11).
@@ 5.1 Task planning and Execution
Task (re-)planning and execution are crucial for the efficient operation of autonomous machines, particularly in dynamic, unstructured environments, enabling them to adapt to changing conditions. In the proposed automated system, task planning and execution are achieved using a behavior tree-based approach. Behavior trees provide a flexible, modular framework for modeling decision-making processes and task execution in a hierarchical, reactive manner [75]. A behavior tree consists of action nodes (which perform tasks), condition nodes (which evaluate specific states), and composite nodes (which control the flow of execution, such as sequences and fallbacks). This structure allows for scalable and adaptable behavior design, making it particularly suited for managing complex tasks like navigation and object manipulation. The integrated behavior tree logic is based on the BehaviorTree.CPP library [76], which offers a performant set of tools for behavior tree construction and execution. To enable smooth communication with the robot’s ROS 2 based architecture, the ROS 2 wrappers from the Nav2 framework [77] are utilized. A simplified version of the behavior tree that describes the operation of ADAPT is shown in Figure 13. The flow of execution can be summarized as follows:
@@
    The ApproachPallet action guides ADAPT to navigate toward the selected pallet.
@@
The full behavior tree is significantly more complex, featuring 26 custom sub-trees and over 30 unique action and composite nodes. It integrates both reactive and anticipatory behaviors. For example, a reactive behavior occurs after a failed pallet pick-up attempt, due to e.g., inaccuracies in the estimated pose, triggering a recovery action and another pick-up attempt.
To initiate the operation, i.e. the execution of the behavior tree, ADAPT receives a command by the forklift supervisor to transport either a specified number of pallets or all available pallets from a designated loading zone to an unloading zone. The current system iteration features a tablet-based HMI for monitoring and controlling the vehicle’s operation. The framework also supports testing of novel interaction modalities, like gesture control, freeing the hands of an operator, see [78] for details.
@@ 5.2.1 Vehicle Base Kinematics
where the Cartesian coordinates of the front axle are defined as $[x_{F},y_{F}]$, with its orientation represented by the angle $\theta_{F}$. The rear part ${\mathbf{q}_{R}}=[x_{R},y_{R},\theta_{R}]^{\mathrm{T}}$ is connected to the front part at the articulation point, with its orientation defined by $\theta_{R}=\pi+\theta_{F}-\gamma$. The kinematics are characterized by the system inputs $v$ and $\dot{\gamma}$ which are the velocity at the front axle and the steering rate, respectively. The parameters $l_{F}$ and $l_{R}$ are the center-to-axle distances from the front as well as the rear axle to the articulation point.
@@ 5.2.2 Hydraulic Modeling
(c) Lift
(d) Shift
Figure 15: Results for modeling the hydraulic drive as well as the hydraulic cylinders showing the measured system output $y_{i}$ compared to the estimated system output $\hat{y_{i}}$, calculated from the actuation command with $i\in\{v,\gamma,l,s\}$. (a)-(b) show data from a path following scenario, where as (c)-(d) show a pallet pickup maneuver.
@@
$\tau_{\zeta}\cdot\dot{\hat{y}}_{\zeta}(t)+\hat{y}_{\zeta}(t)=K_{\zeta}\cdot u_{\zeta}(t)\ ,$ (3)
where the input $u_{\zeta}$ represents the desired valve spool position for the corresponding actuated cylinder and $\hat{y}_{\zeta}$ denotes the estimated output of the system, $K_{\zeta}$ a model constant and $\tau_{\zeta}$ a time constant for the model of $\zeta\in\{\beta,l,s\}$ tilting, lifting and shifting, respectively. The results of the parameter identification experiments are shown in Figure 15, applied to a path following (Figure 15(b)) as well as a pallet pickup scenario (Figure 15(c)-15(d)). The estimated cylinder motion, denoted as $\hat{y}_{\zeta}$, aligns closely with the measured variable $y_{\zeta}$. However, when multiple cylinders are actuated simultaneously, the model’s accuracy decreases due to altered pressure distribution and flow division, as well as pump limitations, leading to non-linear behavior. This reduction in accuracy is not significant in operation, as it can be compensated for by the feedback control system, discussed in the next section.
@@ 5.3 Motion Planning and Control
The operation of ADAPT is divided into two primary action types: (1) Navigation between poses, comprising path planning and following, and (2) Manipulation for loading or unloading of pallets. These primary actions share common vehicle functions, namely vehicle base and fork positioning control. These functions are implemented as two separate cascaded control structures, as illustrated in Figure 16. The main components of these control loops are the pose tracking control (PTC) and fork tip transformation (FTT) which track the external reference. The inner control loop consists of a feed-forward control (FF) based on the hydraulic models from Section 5.2.2 as well as basic PI(D) control and saturation blocks for hydraulic valve and cylinder control. This section delves into the details of the primary action types and explains how vehicle functions are implemented.
@@ 5.3.1 Cascaded control
Figure 16: Cascaded control architectures for the vehicle base (a) and fork actuation (b), forming the foundation for pose-to-pose navigation and pallet manipulation. Main components, enclosed in red boxes, are the pose tracking control (PTC) and fork tip transformation (FTT) and run on the IPC, while those in cyan boxes operate on the real-time PLC. For navigation, FTT is not needed.
@@
which is calculated by transforming the Cartesian coordinates of the reference vehicle into the vehicle’s body-fixed origin ${\mathbf{q}_{F/R}}=[x_{F/R},y_{F/R},\theta_{F/R}]^{\mathrm{T}}$. This can either be the center of the front (F) or the rear (R) axle. The state comprises the relative errors in longitudinal direction $e_{x}$, lateral direction $e_{y}$ and of the heading $e_{\theta}$. The approach neglects $e_{x}$ and instead focuses on minimizing $e_{y}$ and $e_{\theta}$ which results in the control law for the steering rate
@@
which is designed to respect the constrained system’s steering dynamics. In here, $v_{R}$ is the desired reference velocity for operation. The parameter $k_{v}$ allows to reduce the velocity proportional to the squared desired steering rate. Thus, sharp curvatures are tracked with lower speed, to maintain tracking accuracy. The minimum velocity $v_{\min}$ ensures that the vehicle maintains a minimum speed, even during sharp steering maneuvers. Equations (4)-(6) build the foundation of the PTC and, thus, the cascaded control loop in Figure 16a.
Fork positioning control: For a successful pallet manipulation, an accurate positioning of the fork tip is necessary. The most critical is pallet pickup, where the tolerated positioning error must not exceed 5 cm. To this end, the tip of the fork $\mathbf{q}_{T}=[x_{T},y_{T},\theta_{T},z_{T}]^{\mathrm{T}}$ must be precisely aligned with the pallet front $\mathbf{q}_{P}=[x_{P},y_{P},\theta_{P},z_{P}]^{\mathrm{T}}$, see Figure 19(b). This is ensured by the fork tip transformation (FTT) component, as part of the cascaded control loop depicted in Figure 16b. The FTT transforms the target pose into the fork tip coordinate frame by
@@
Based on (7), the reference signal for the fork shift corresponds to the lateral error, $s^{d}=y_{P}^{\prime}$, and the lift mast height to the vertical error, $l^{d}=z_{P}$. The tilting command $\beta^{d}$ is adjusted to keep the forks parallel to the ground. The cascaded loop in Figure 16b can operate without the FTT, depending on the forklift’s current task. Specifically, during transporting a pallet - while not loading or unloading - the FTT becomes optional, allowing the setpoints for the underlying control loop $l^{d},s^{d}$ and $\beta^{d}$ to be directly assigned to a transport position that keeps the fork at a safe height and angle.
@@ 5.3.2 Navigation
The Navigation action comprises the planning and subsequent following of a collision free path from one pose to another. Planning a path is a well-established research area for robotic applications, with numerous feasible algorithms. Empirical evaluation of potential path planners led to the conclusion that the Hybrid A* stands out as particularly advantageous for flexible path planning in construction environments. This is due to its ability to plan near-optimal bi-directional paths while maintaining computational efficiency, and it places no restrictions on path length or the number and location of changes of driving direction. Additionally, the preferred driving direction can be forced with an adjustable penalty term.
However, the Hybrid A* algorithm is formulated for car-like vehicles and cannot be directly applied to articulated vehicles. Nonetheless, for symmetrical articulated vehicles, i.e. $l_{R}=l_{F}$, the kinematic properties for static curvature and straight-line driving are equal to those of car-like vehicles. This is achieved by superimposing the Instantaneous Centre of Rotation (ICR), i.e. the intersection point of the front and rear axles, of both vehicle models.
Geometric transformations relate the steering angle of the articulated vehicle
$\gamma=\arctan{\left(\sin\left(\frac{\gamma_{cl}}{2}\right)\sqrt{2\left(\cos\left(\alpha\right)\cos\left(\beta\right)+1\right)}\right)}$ (8)
to the steering angle of the car-like model $\gamma_{cl}$. The relation between steering angle and articulation angle $\alpha$ follows from the rotation between the front and rear part
@@
These two equations (8) and (9) can be solved numerically for the steering angle $\gamma$ or the articulation angle $\alpha$. Figure 17 shows the deviation between the planned angle $\gamma_{cl}$ for car-like steering and the kinematic-matching articulation angle $\alpha$.
@@
Since the forklift is designed to operate in both forward and reverse driving directions and the implemented path tracking control law is designed for uni-directional movement, the bi-directional path following approach employs a multi-stage approach. The bi-directional path consists of segments, separated by turning poses, with alternating movement direction. Therefore, the respective vehicle pose also alternates between the front $\mathbf{q}_{F}$ and the rear axle $\mathbf{q}_{R}$. This approach is facilitated by the forklift’s symmetric geometry and smooth transitions between the segments are handled by the behavior tree. An example of bi-directional path following, including a turning pose, is shown in Figure 18.
In addition to controlling the vehicle base, the path-following module utilizes the simplified fork control loop, see Figure 16b, without FTT) to ensure that the forks maintain a safe height and angle, particularly when transporting a load. This is achieved by directly applying the predefined setpoint values $l^{d},s^{d}$ and $\beta^{d}$ for the respective joints.

### v2

@@ 4.4 Loading Edge Detection
Outlier removal: Outliers are removed by discarding points with insufficient number of neighbors within a specified radius.
@@
Figure 11: A LiDAR point cloud (colored by reflectance) captures the scene, including the forklift forks (bottom right) and their shadow. The lift is positioned at an estimated loading height to optimize the platform’s visibility. The resulting detected loading edge reference frame (RGB frame) and the target slots (red boxes) defined within this frame are depicted as overlay.
@@ 4.5 Obstacle Detection
(b) Birds-eye view: Potential obstacles (green) are tracked to assess movement but are only considered as future collisions (red) if they intersect with the path volume (cyan).
@@
Figure 12: Depiction of the obstacle detection system’s geometric relations and output. Definitive collision objects (red) that trigger emergency braking overlap with the expected path volume (cyan), and tracked potential obstacles are outside (green).
@@
For safe operation, the forklift must detect obstacles quickly to avoid collisions with static obstacles (trees, walls, etc.) and moving obstacles (excavators, pedestrians, etc.), which may approach the currently planned drive path. The implemented approach is a fast-reacting method in the sensor frame complementary to and independent of the mapping approach presented in Section 4.2. It is applied only in the forward driving direction, with a higher risk of colliding with moving obstacles. In contrast, movements in the fork direction, as discussed in Section 3.3.2, are slower and follow much shorter paths. Based on the selected safety concept, when an obstacle is detected, the forklift is brought to a safe state (standstill) to prevent collisions, rather than re-planning its path to navigate around the obstacle. Once an emergency stop is triggered, the forklift remains stationary until it receives operator input to resume the task.
@@
For the assessment of the risk of collision with moving obstacles, we utilize a multi-object detection and tracking system, originally based on the works [72] and [73], which further became the basis for an endurance-tested tramway braking assistance system in commercial use [^5] https://www.alstom.com/solutions/rolling-stock/citadis-light-rail-designed-reflect-your-citys-unique-identity (accessed 2024-12-11). It operates in a tracking-by-detection paradigm. It can detect arbitrary objects, track them over time, and assess if a collision will occur based on the predicted trajectories of the objects and the projected path of the forklift. The detection step is based on generic 3D clustering. Hence, it does not depend on predefined object classes such as persons or vehicles.
@@
In order to detect obstacles, the system analyzes point clouds from the Ouster LiDAR mounted on the front part of the forklift. Both LiDAR and a custom stereo camera setup, see Figure 3c, have been successfully used, but the wider field of view of the LiDAR system makes it the safer choice for obstacle detection, see Figure 12. Points above an assumed ground plane, cf. Figure 12(a), are analyzed for potential obstacles by grouping them into separate distinct objects using Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [74]. The clusters represent potential obstacles (indicated as red and green boxes in Figure 12). However, only if they overlap with the expected vehicle’s path are they considered potential collision objects (red) that trigger an emergency brake. The volume of the monitored path, cyan in Figure 12, is defined by extruding the vehicle footprint (plus an additional clearance buffer) along a simplified vehicle path. Obstacle bookkeeping is implemented similarly to pallet bookkeeping, as described in Section 4.1. However, it is enhanced with a Kalman filter based on a constant velocity model to accommodate movement. Then each object track is checked for a possible collision, which can cause a collision warning based on the predicted movement of the object, the velocity of the forklift, the characteristics of the forklift deceleration, and a defined minimum safety distance.
@@ 5 Planning and Control
ADAPT performs two primary types of actions to carry out its tasks effectively: (1) collision-free navigation, where the forklift moves from its current position to a target location within a predefined operational area, either while carrying a pallet or traveling unloaded, and (2) precise manipulation, which involves the accurate pick-up and placement of pallets in specified locations (slots), such as on a truck or within a designated area on the ground (multiple slots on a truck are shown in Figure 11).
@@ 5.1 Task Planning and Execution
Task (re-)planning and execution are crucial for the efficient operation of autonomous machines, particularly in dynamic, unstructured environments, enabling them to adapt to changing conditions. In the proposed autonomous system, task planning and execution are achieved using a behavior tree-based approach. Behavior trees provide a flexible, modular framework for modeling decision-making processes and task execution in a hierarchical, reactive manner [75]. A behavior tree consists of action nodes (which perform tasks), condition nodes (which evaluate specific states), and composite nodes (which control the flow of execution, such as sequences and fallbacks). This structure allows for scalable and adaptable behavior design, particularly suited for managing complex tasks like navigation and object manipulation. The integrated behavior tree logic is based on the BehaviorTree.CPP library [76], which offers a performant set of tools for behavior tree construction and execution. In order to enable smooth communication with the robot’s ROS 2 based architecture, the ROS 2 wrappers from the Nav2 framework [77] are utilized. A simplified version of the behavior tree that describes the operation of ADAPT is shown in Figure 13. The flow of execution can be summarized as follows:
@@
    The ApproachPallet action guides ADAPT in navigating toward the selected pallet.
@@
The full behavior tree is significantly more complex, featuring 26 custom sub-trees and over 30 unique action and composite nodes. It integrates both reactive and anticipatory behaviors. For example, a reactive behavior occurs after a failed pallet pick-up attempt, due to, for example, inaccuracies in the estimated pose, triggering a recovery action and another pick-up attempt.
To initiate the operation, i.e., the execution of the behavior tree, ADAPT receives a command by the forklift supervisor to transport either a specified number of pallets or all available pallets from a designated loading zone to an unloading zone. The current system iteration features a tablet-based HMI for monitoring and controlling the vehicle’s operation. The framework also supports the testing of novel interaction modalities, like gesture control for freeing the hands of an operator; see [78] for details.
@@ 5.2.1 Vehicle Base Kinematics
where the Cartesian coordinates of the front axle are defined as $[x_{F},y_{F}]$, with its orientation represented by the angle $\theta_{F}$. The rear part ${\mathbf{q}_{R}}=[x_{R},y_{R},\theta_{R}]^{\mathrm{T}}$ is connected to the front part at the articulation point, with its orientation defined by $\theta_{R}=\pi+\theta_{F}-\gamma$. The kinematics are characterized by the system inputs $v$ and $\dot{\gamma}$, which are the velocity at the front axle and the steering rate, respectively. The parameters $l_{F}$ and $l_{R}$ are the center-to-axle distances from the front and the rear axle to the articulation point.
@@ 5.2.2 Hydraulic Modeling
(c) Tilt
(d) Lift
(e) Shift
Figure 15: Results for modeling the hydraulic drive and the hydraulic cylinders showing the measured system output $y_{i}$ compared to the estimated system output $\hat{y_{i}}$, calculated from the actuation command with $i\in\{v,\gamma,\beta,l,s\}$. Subfigures (a) and (b) show data from a path following scenario, and Subfigures (c)-(e) refer to a pallet pick-up maneuver.
@@
$\tau_{\zeta}\dot{\hat{y}}_{\zeta}(t)+\hat{y}_{\zeta}(t)=K_{\zeta}u_{\zeta}(t)\ ,$ (3)
where the input $u_{\zeta}$ represents the desired valve spool position for the corresponding actuated cylinder and $\hat{y}_{\zeta}$ denotes the estimated output of the system, $K_{\zeta}$ a model constant, and $\tau_{\zeta}$ a time constant for the model of $\zeta\in\{\beta,l,s\}$ tilting, lifting, and shifting, respectively. The results of the parameter identification experiments are shown in Figure 15, applied to a path following (Figure 15(b)) and a pallet pick-up scenario (Figure 15(c)-15(e)). The estimated cylinder motion, denoted as $\hat{y}_{\zeta}$, aligns closely with the measured variable $y_{\zeta}$. However, when multiple cylinders are actuated simultaneously, the model’s accuracy decreases due to altered pressure distribution, flow division, and pump limitations, leading to nonlinear behavior. This reduction in accuracy is not significant in operation, as it can be compensated for by the feedback control system, discussed in the next section.
@@ 5.3 Motion Planning and Control
The operation of ADAPT is divided into two primary action types: (1) Navigation between poses, comprising path planning and following, and (2) manipulation for loading or unloading pallets. These primary actions share common vehicle functions, i.e., vehicle base and fork positioning control. These functions are implemented as two separate cascaded control structures, as illustrated in Figure 16. The main components of these control loops are the pose tracking control (PTC) and fork tip transformation (FTT), which track the external reference. The inner control loop consists of a feedforward control (FF) based on the hydraulic models from Section 5.2.2 and basic PI(D) control and saturation blocks for hydraulic valve and cylinder control. This section delves into the details of the primary action types and explains how vehicle functions are implemented.
@@ 5.3.1 Cascaded control
Figure 16: Cascaded control architectures for the vehicle base (a) and fork actuation (b) form the foundation for pose-to-pose navigation and pallet manipulation. The main components, enclosed in red boxes, are the pose tracking control (PTC) and fork tip transformation (FTT), which run on the IPC, while those in cyan boxes operate on the real-time PLC. For navigation, FTT is not needed.
@@
which is calculated by transforming the Cartesian coordinates of the reference vehicle into the vehicle’s body-fixed origin ${\mathbf{q}_{F/R}}=[x_{F/R},y_{F/R},\theta_{F/R}]^{\mathrm{T}}$. This can either be the center of the front (F) or the rear (R) axle. The state comprises the relative errors in longitudinal direction $e_{x}$, lateral direction $e_{y}$, and the heading $e_{\theta}$. The approach neglects $e_{x}$ and instead focuses on minimizing $e_{y}$ and $e_{\theta}$ which results in the control law for the steering rate
@@
which is designed to respect the constrained system’s steering dynamics. Here, $v_{R}$ is the desired reference velocity for operation. The parameter $k_{v}$ allows the reduction of the velocity proportional to the squared desired steering rate. Thus, sharp curvatures are tracked at a lower speed to maintain tracking accuracy. The minimum velocity $v_{\min}$ ensures that the vehicle maintains a minimum speed, even during sharp steering maneuvers. Equations (4)-(6) build the foundation of the PTC and, thus, the cascaded control loop in Figure 16a.
Fork positioning control: For a successful pallet manipulation, an accurate positioning of the fork tip is necessary. The most critical is pallet pick-up, where the tolerated positioning error in lateral and vertical direction must not exceed 5 cm. To this end, the tip of the fork $\mathbf{q}_{T}=[x_{T},y_{T},\theta_{T},z_{T}]^{\mathrm{T}}$ must be precisely aligned with the pallet front $\mathbf{q}_{P}=[x_{P},y_{P},\theta_{P},z_{P}]^{\mathrm{T}}$, see Figure 20(b). This is ensured by the fork tip transformation (FTT) component, as part of the cascaded control loop depicted in Figure 16b. The FTT transforms the target pose into the fork tip coordinate frame by
@@
Based on (7), the reference signal for the fork shift corresponds to the lateral error, $s^{d}=y_{P}^{\prime}$, and the lift mast height to the vertical error, $l^{d}=z_{P}$. The tilting command $\beta^{d}$ is adjusted to keep the forks parallel to the ground. The control loop relies on the vehicle base control, which minimizes the orientation error $e_{\theta}$. If $e_{\theta}$ or $y^{\prime}_{P}$ exceed a predetermined threshold shortly before the forks enter the pallet (i.e., $\epsilon_{x}>x^{\prime}_{P}>0$, $\epsilon_{x}=0.2$m), the pallet pick-up procedure is aborted, and a recovery maneuver is initiated via the task planner.
The cascaded loop in Figure 16b can operate without the FTT, depending on the forklift’s current task. Specifically, during transporting a pallet - while not loading or unloading - the FTT becomes optional, allowing the setpoints for the underlying control loop $l^{d},s^{d}$, and $\beta^{d}$ to be directly assigned to a transport position that keeps the fork at a safe height and angle.
@@ 5.3.2 Navigation
The navigation action comprises planning and following a collision-free path from one pose to another. Planning a path is a well-established research area for robotic applications, with numerous feasible algorithms. Based on an empirical evaluation of potential path planners, the Hybrid A* is advantageous for flexible path planning in construction environments. This is due to its ability to plan near-optimal bi-directional paths while maintaining computational efficiency, and it places no restrictions on path length or the number and location of driving direction changes. Moreover, an adjustable penalty term can be employed to prefer specific driving directions.
However, the Hybrid A* algorithm is formulated for car-like vehicles and cannot be directly applied to articulated vehicles. Nonetheless, for symmetrical articulated vehicles, i.e., $l_{R}=l_{F}$, the kinematic properties for static curvature and straight-line driving are equal to those of car-like vehicles. This is achieved by superimposing the Instantaneous Centre of Rotation (ICR), i.e., the intersection point of the front and rear axles, of both vehicle models. Fig. 17 illustrates the articulated vehicle model with its ICR, the articulation angle $\alpha$, and the tilting angle $\beta$.
[…]
Figure 17: ICR for the articulated vehicle in a tilted position: Horizontal shading marks the angle in the horizontal plane and vertical shading in the vertical plane.
Using the geometric relations of the front part, we obtain the steering angle of the articulated vehicle
$\gamma=\arctan{\left(\frac{l_{R}}{2r_{v}}\sqrt{2\left(\cos\left(\alpha\right)\cos\left(\beta\right)+1\right)}\right)}$ (8)
with the radius of the ICR $r_{v}$. For the car-like model, the same radius of the ICR results in
$r_{v}=\frac{l_{R}}{2\sin\left(\gamma_{cl}/2\right)},$ (9)
with the steering angle of the car-like model $\gamma_{cl}$. The relation between the steering angle and the articulation angle $\alpha$ follows from the rotation between the front and rear part
@@
These two equations (8) and (10), together with (9), can be solved numerically for the steering angle $\gamma$ or the articulation angle $\alpha$. Figure 18 shows the deviation between the planned angle $\gamma_{cl}$ for car-like steering and the kinematic-matching articulation angle $\alpha$.
@@
Since the forklift is designed to operate in both forward and reverse driving directions and the implemented path tracking control law is designed for uni-directional movement, the bi-directional path following approach employs a multi-stage approach. The bi-directional path consists of segments, separated by turning poses, with alternating movement direction. Therefore, the respective vehicle pose also alternates between the front $\mathbf{q}_{F}$ and the rear axle $\mathbf{q}_{R}$. This approach is facilitated by the forklift’s symmetric geometry and the behavior tree handles smooth transitions between the segments. An example of bi-directional path following, including a turning pose, is shown in Figure 19.
In addition to controlling the vehicle base, the path-following module utilizes the simplified fork control loop, see Figure 16b, without FTT) to ensure that the forks maintain a safe height and angle, particularly when transporting a load. This is achieved by directly applying the predefined setpoint values $l^{d},s^{d}$, and $\beta^{d}$ for the respective joints.